OpenStudy (anonymous):

a + bi + 0 + 0i = a + bi

OpenStudy (anonymous):

I'd say this is a true statement. 0 + 0i = 0 so a +bi=a+bi

OpenStudy (anonymous):

a+bi+0+0i=a+bi a+bi=a+bi.